OpenStudy (anonymous):

the are of a rectangular swimming pool is given by 4x^2+19x+12 ft^2, one side length of the pool is given by 4x+3 feet. What is an algebraic expression for the other side of length of pool? Simplify this.

OpenStudy (lgbasallote):

remember that area = length x width

OpenStudy (anonymous):

as area is equal to length times height, you can make an equation as follows: \[(4x+3) \times (y) = (4x^2 + 19 + 12)\] where y is the other side. next factor the equation \[(4x^2 + 19 + 12)\]

OpenStudy (anonymous):

can you do it from there? or do you need me to explain more?

OpenStudy (anonymous):

I need you explain if you do not mind. I am new to this. Very new, first time at college at an older age and really trying to learn but I get confused. Would appreciate the help!!

OpenStudy (anonymous):

that's fine :) ok so as we know that to factor \[( 4x^2 + 19x +12) \] one of the solutions is going to be \[(4x +3) \] from looking at the equation we can factor the larger equation easily :) so: \[( 4x^2 + 19x +12) = (4x + 3)(x+4) \] do you understand the factoring?

OpenStudy (anonymous):

Yes, i understand factoring

OpenStudy (anonymous):

ok great well from there we find that the other side (y as I mentioned earlier) is equal to (x +4)

OpenStudy (anonymous):

ok. these numbers get really complicated!!

OpenStudy (lgbasallote):

great teaching @bbsmilieface ;D keep it up! ^_^

OpenStudy (anonymous):

@lgbasallote thanks :) @chicken2 I hope you understand now :)

OpenStudy (anonymous):

I will continue to work on some other practice problems, but with the way yoou explained it, it really helped. Thank you so much. Like I said ealier all of this is new to me.